# DeepCompress: A Dual Reward Strategy for Dynamically Exploring and Compressing Reasoning Chains

### v1

@@ 1 Introduction
In this paper, we propose DeepCompress, a novel framework that incorporates an adaptive length reward mechanism which dynamically adjusts the preference for shorter or longer responses based on the problem difficulty perceived in real-time by the LRMs. Specifically, we first reveal that longer responses contain a wider coverage of potentially correct solutions than shorter ones for the same problems. In other words, current methods that constantly optimize shorter responses in RL processes may constrain the problem-solving capacity of LRMs and restrict their reasoning boundary. At the meantime, it is infeasbile to encourage LRMs to always favor longer responses in RL processes considering the effciency for both training and inference. Our DeepCompress addresses this challenge by first dividing the problems into “Simple“ or “Hard“ classes and then applying different length reward modes to them, respectively, during training. With Group Relative Policy Optimization (GRPO, Shao et al., 2024) as our basic RL algorithm, we consider a problem as “Simple“ when its group pass ratio (i.e., the proportion of correct samples among its $G$ generated responses) exceeds the batch pass ratio (i.e., average of group pass ratio in the batch), and “Hard“ when otherwise. Then, DeepCompress encourages the LRMs to favor shorter responses of the “Simple“ problems, but longer responses of the “Hard“ problems. Through this mechanism, DeepCompress dynamically adapts the reasoning chain length – autonomously compressing lengthy CoT for well-mastered problems while extending CoT for under-learned cases.
@@
  • We propose DeepCompress, which incorporates a model-aware diffifulty mechanism to dynamically classify questions as “Simple” or “Hard”, and a dual length reward to adaptively explore longer responses for “Hard” questions and favor shorter responses for “Easy” ones.
@@ 3.1 Experimental Setup
To assess the mathematical performance of our models, we followed Zeng et al. (2025) and evaluated them on four challenging benchmarks: MATH-500 (Hendrycks et al., 2021), OlympiadBench (He et al., 2024), Minerva Math (Lewkowycz et al., 2022) and AIME 2025 (MAA, a).
@@ 4.3 Model-Aware Difficulty
A core challenge in implementing the dual length reward is the dynamic assessment of question difficulty. While public datasets like MATH (Hendrycks et al., 2021) and DeepMath (He et al., 2025) offer curated difficulty labels, this approach incurs additional annotation costs and fails to adapt to the model’s evolving capabilities during training.
@@ 5.1 Experimental Setup
We comprehensively evaluate model performance on seven challenging mathematical benchmarks: MATH-500 (Hendrycks et al., 2021), AMC 2023 (MAA, b), OlympiadBench (He et al., 2024), Minerva Math (Lewkowycz et al., 2022), AIME 2024-2025 (MAA, a), and the English subset of PolyMath (Wang et al., 2025a). As primary metrics, we samples 16 responses for each question and report the pass@1 accuracy. We construct a validation set, which consists of 60 questions from MATH and 60 from AIME 2022-2023, to select the checkpoint with the highest pass@1 score for evaluation. We utilized vLLM (Kwon et al., 2023) for efficient batch inference, and fixed the decoding parameters to temperature=0.6, top_p=0.95, and max_tokens=32,768. To ensure fair comparison and eliminate variance from evaluation scripts, we re-evaluate all baseline models under our precise evaluation settings.
Table 1: Math reasoning performance. “DeepCompress” denotes models trained with our novel DeepCompress approach, which improves the reasoning accuracy and efficiency simultaneously.
@@ 5.4 Quantifying Emergence of Reasoning Behaviors
To further investigate the mechanisms behind DeepCompress’s high policy entropy and superior performance, we conducted an analysis on a targeted set of challenging problems. Specifically, we constructed this hard problem set from instances where our baseline models (Qwen2.5-3B and Qwen2.5-7B), failed to produce a correct solution. On each set, we follow Zeng et al. (2025) to track the emergence of four cognitive behaviors described in Gandhi et al. (2025). The manifestation of these behaviors suggests a reproduction of the “aha moment” phenomenon observed in R1 (Guo et al., 2025).
Table 2: Reflection Frequency on hard questions. We use GPT-4o to extract and track “aha moment” behaviors, with the prompt shown in Appendix A.
@@ 6 Conclusion
Appendix A Reasoning Behavior Prompt
@@ Appendix B Training Details
total_training_steps 500 500

### v2

@@ 1 Introduction
In this paper, we propose DeepCompress, a novel framework that incorporates an adaptive length reward mechanism which dynamically adjusts the preference for shorter or longer responses based on the problem difficulty perceived in real-time by the LRMs. Specifically, we first reveal that longer responses contain a wider coverage of potentially correct solutions than shorter ones for the same problems. In other words, current methods that constantly optimize shorter responses in RL processes may constrain the problem-solving capacity of LRMs and restrict their reasoning boundary. At the meantime, it is infeasible to encourage LRMs to always favor longer responses in RL processes considering the efficiency for both training and inference. Our DeepCompress addresses this challenge by first dividing the problems into “Simple“ or “Hard“ classes and then applying different length reward modes to them, respectively, during training. With Group Relative Policy Optimization (GRPO, Shao et al., 2024) as our basic RL algorithm, we consider a problem as “Simple“ when its group pass ratio (i.e., the proportion of correct samples among its $G$ generated responses) exceeds the batch pass ratio (i.e., average of group pass ratio in the batch), and “Hard“ when otherwise. Then, DeepCompress encourages the LRMs to favor shorter responses of the “Simple“ problems, but longer responses of the “Hard“ problems. Through this mechanism, DeepCompress dynamically adapts the reasoning chain length – autonomously compressing lengthy CoT for well-mastered problems while extending CoT for under-learned cases.
@@
  • We propose DeepCompress, which incorporates a model-aware mechanism to dynamically classify questions by difficulty, and a dual length reward to adaptively explore longer responses for “Hard” questions and favor shorter responses for “Simple” ones.
@@ 3.1 Experimental Setup
To assess the mathematical performance of our models, we followed Zeng et al. (2025) and evaluated them on four challenging benchmarks: MATH-500 (Hendrycks et al., 2021b), OlympiadBench (He et al., 2024), Minerva Math (Lewkowycz et al., 2022) and AIME 2025 (MAA, a).
@@ 4.3 Model-Aware Difficulty
A core challenge in implementing the dual length reward is the dynamic assessment of question difficulty. While public datasets like MATH (Hendrycks et al., 2021b) and DeepMath (He et al., 2025) offer curated difficulty labels, this approach incurs additional annotation costs and fails to adapt to the model’s evolving capabilities during training.
@@ 5.1 Experimental Setup
We comprehensively evaluate model performance on seven challenging mathematical benchmarks: MATH-500 (Hendrycks et al., 2021b), AMC 2023 (MAA, b), OlympiadBench (He et al., 2024), Minerva Math (Lewkowycz et al., 2022), AIME 2024-2025 (MAA, a), and the English subset of PolyMath (Wang et al., 2025a). As primary metrics, we samples 16 responses for each question and report the pass@1 accuracy. We construct a validation set, which consists of 60 questions from MATH and 60 from AIME 2022-2023, to select the checkpoint with the highest pass@1 score for evaluation. We utilized vLLM (Kwon et al., 2023) for efficient batch inference, and fixed the decoding parameters to temperature=0.6, top_p=0.95, and max_tokens=32,768. To ensure fair comparison and eliminate variance from evaluation scripts, we re-evaluate all baseline models under our precise evaluation settings.
Table 1: Math reasoning performance. “DeepCompress” denotes models trained with our novel DeepCompress approach, which significantly improves the reasoning accuracy.
@@ 5.2 Main Results
Table 2: Performance on the GPQA-Diamond, MMLU-STEM and Big-Bench Hard.
Model GPQA-Diamond MMLU Big-Bench
Biology Physics Chemistry Overall STEM Hard
Qwen-2.5-3B 29.9 19.8 20.3 21.0 7.5 48.1
Qwen-2.5-3B-Instruct 45.1 26.1 30.7 29.9 48.7 71.3
DeepMath-Zero-3B 45.1 25.3 32.2 30.2 54.7 71.6
DeepCompress-Zero-3B 44.1 25.5 35.7 31.7 56.0 73.7
Qwen-2.5-7B 33.6 21.4 27.8 25.3 12.1 41.5
Open-Reasoner-Zero-7B 50.3 26.7 47.8 38.1 47.0 83.2
Qwen-2.5-7B-SimpleRL-Zoo 31.9 22.6 37.9 30.2 15.0 74.9
DeepMath-Zero-7B 58.6 29.5 53.2 42.6 72.7 85.0
DeepCompress-Zero-7B 57.6 31.2 58.2 43.9 75.5 85.7
Generalizable reasoning beyond mathematics
To validate that our method generalizes beyond mathematical reasoning, we added evaluations on GPQA-Diamond (Rein et al., 2024) (biology, physics and chemistry), Big Bench Hard (Hendrycks et al., 2021a) (complex multi-step reasoning), and MMLU-STEM (Suzgun et al., 2023) (science and engineering reasoning). As shown in Table 2, DeepCompress models consistently outperform other baselines, demonstrating significant improvements across all three benchmarks. These findings suggest that the dynamic strategy of exploration and efficiency facilitated by DeepCompress does not merely optimize for mathematical tasks, but rather fosters a more robust reasoning capability that generalizes across diverse scientific domains.
@@ 5.4 Quantifying Emergence of Reasoning Behaviors
To further investigate the mechanisms behind DeepCompress’s high policy entropy and superior performance, we conducted an analysis on a targeted set of challenging problems. Specifically, we constructed this hard problem set from instances where our baseline models (Qwen2.5-3B and Qwen2.5-7B), failed to produce a correct solution. On each set, we follow Zeng et al. (2025) to track the emergence of four cognitive behaviors described in Gandhi et al. (2025), with the prompt shown in Appendix D. The manifestation of these behaviors suggests a reproduction of the “aha moment” phenomenon observed in R1 (Guo et al., 2025).
Table 3: Reflection Frequency on hard questions. We use GPT-4o to extract and track “aha moment” behaviors, with the prompt shown in Appendix D.
@@
5.5 Hyperparameter Analysis
Figure 5: Ablation studies on $\alpha$ and $\lambda$. The dashed line represents the DeepMath-Zero-3B baseline.
Reward Weight $\alpha$ and EMA Parameter $\lambda$
As illustrated in Figure 5, our method exhibits remarkable robustness with respect to the reward weight $\alpha$, with all configurations consistently outperforming the previous baseline, DeepMath-Zero-3B. Additionally, empirical results indicate that a larger smoothing factor ($\lambda=0.99$) is crucial for achieving optimal performance. By incentivizing longer reasoning trajectories and deeper exploration during the early training phase, this elevated factor effectively empowers the model to break through existing performance bottlenecks.
@@ 6 Conclusion
7 Ethics Statement
The authors of this work have read and adhere to the ICLR Code of Ethics. Our research focuses on developing a novel reinforcement learning framework, DeepCompress, aimed at enhancing the reasoning capabilities and computational efficiency of Large Reasoning Models. The primary goal of our work is to advance the scientific understanding of AI reasoning and to develop models that are both more effective and more efficient, which we believe is a positive step towards sustainable and accessible AI research. Our work is foundational and does not introduce societal harms. All code and models will be released publicly to promote open research and reproducibility.
8 Reproducibility Statement
To ensure full reproducibility, all our code, training scripts, and final model weights will be made publicly available. Detailed descriptions of our training setup, including all hyperparameters, software versions, and implementation specifics, are provided in Appendix B. This will allow researchers to verify our results, build upon our framework, and further explore adaptive training strategies for large reasoning models.
Appendix A The Use of Large Language Models
In the preparation of this manuscript, we utilized a Large Language Model as a writing assistant. Its role was to aid in polishing the text, including improving grammar and sentence structure. The core research ideas, experimental design, and analysis presented in this paper are entirely our own.
@@ Appendix B Training Details
total_training_steps 600 600
@@
Appendix C Batch Pass Ratio
The core mechanism of DeepCompress relies on the batch pass ratio ($P_{b}$) to judge problem difficulty. We recorded the changes in $P_{b}$ throughout the training process, and as shown in Figure 6, the model’s $P_{b}$ exhibits stable growth, indicating very low noise in the difficulty judgment process.
Figure 6: Batch pass ratio during training.
Appendix D Reasoning Behavior Prompt